On the day his grandson was born, a man deposited with a trust company a sufficient amount of money so that the boy could receive five annual payments of P2,000 each starting with his 18th birthday. Interest at the rate of 9% per annum was to be paid on all amounts on deposit. There also was a provision that the grandson could elect to withdraw no annual payments and receive a sing lump amount on his 23rd birthday. The grandson chose this option. How much did the boy receive as the single payment?


1
Expert's answer
2021-11-18T10:21:11-0500

Solution:

Let single payment = A

A = P(1 + r)n

Where: A = End amount plus interest =?

            P = Annual payments = 2,000

            r = Interest rate = 12%

            n = Years = 5yrs

A = 2,000(1 + 0.12)5 = 2,000 x 1.76234 = 3,524.68

 

The boy received on his 23rd birthday a single payment = 3,524.68